OpenStudy (anonymous):

Multiply x^1/4 x x^2/3 x^2/7 x^1/6 x^3/7 x^11/12

jimthompson5910 (jim_thompson5910):

Add the exponents to get 1/4+2/3 = 3/12+8/12 = 11/12 So the final exponent is 11/12 So the answer is D

OpenStudy (anonymous):

x^ (1/4 + 2/3) = x^(11/12)